A test consists of 10 multiple choice questions, each with five possible answers, only one of which is correct. Find the mean and the standard deviation of the number of correct answers.
Answer by Boreal(15235)

mean is np=10*0.2=2 answers
variance is np(1-p)=2*0.8=1.6 answers^2
sd is sqrt(V)=1.26 answers